A problem of statistics is given to 3 students A,B & C whose chances of solving it are 1/2,1/3 & 1/4 respectively.what is the probabilty that the problem can be solved?
1
Expert's answer
2011-10-06T12:33:50-0400
The probability that the problem is not solved by any of them: (1 - P(A))*(1 - P(B))* (1 - P(C)) = (1 - 1/2)(1 - 3/4)(1 - 1/4) = 1/2*1/4*3/4 = 3/32
Hence, the probability that the problem will be solved is 1 - 3/32 = 29/32.